OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined. @dinmosiaren @bibby @adrynicoleb @Evictu_FB @Rockiee @Mikhael

OpenStudy (anonymous):

@KattGirl14 @april115 @Whiteboy1949

OpenStudy (anonymous):

@messiahcollins

OpenStudy (anonymous):

5x + 6 = 41 First, we'll subtract 6 from both sides: 5x + 6 - 6 = 41 - 6 5x = 35 Now divide both sides by 5 5x/5 = 35/5 x = 7 We can check our answer by plugging x = 7 back into the original equation: 5x + 6 = 41 5(7) + 6 = 41 35 + 6 = 41 41 = 41 5x+6 = 41 5x = 41-6 5x = 35 x = 35/5 x = 7

OpenStudy (anonymous):

Simplifying 28 = -2x + 13 Reorder the terms: 28 = 13 + -2x Solving 28 = 13 + -2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x' to each side of the equation. 28 + 2x = 13 + -2x + 2x Combine like terms: -2x + 2x = 0 28 + 2x = 13 + 0 28 + 2x = 13 Add '-28' to each side of the equation. 28 + -28 + 2x = 13 + -28 Combine like terms: 28 + -28 = 0 0 + 2x = 13 + -28 2x = 13 + -28 Combine like terms: 13 + -28 = -15 2x = -15 Divide each side by '2'. x = -7.5 Simplifying x = -7.5